OpenStudy (chase):

(1+tan)(tan2x)=2tanx/1-tanx

OpenStudy (anonymous):

is it : 1+tan(x)?

OpenStudy (anonymous):

after simplifying you'll have 1=1... what do you need to do...?

OpenStudy (chase):

both sides have to be equal to each other

OpenStudy (chase):

how did you simplify it to 1=1?

OpenStudy (anonymous):

tan2x=2tanx/(1-tanx^2) - double-angle formula so, (1+tanx)*2*tanx/(1-tanx^2) = 2tanx/(1-tanx) simplify - 2tans=canceled out on both sides (1+tanx)/[(1-tanx)*(1+tanx)]= 1/(1-tanx) everything canceling out.... so, in general left side of your original eq = right side unless, I mis-read your original problem, it should be right...